# Galactic bars are already mature at Cosmic Noon: bar strength and flatness at $z\sim 1.5$

### v1

@@ 3.2 Masking the spiral arms
3.5 The bar length uncertainty and its implication
We fit the bars in our sample galaxies with a Sérsic profile. However, this does not allow us to measure bar length. Physically, the bar length is the maximum extent of the $x_{\mathrm{1}}$ orbits of the trapped stars making up the bar (binney08). In the local Universe, this typically corresponds to the major axis derived from ellipse fits to surface brightness isophotes (Sec. 3.1), beyond which the bar surface brightness drops rapidly. For this reason, 2D Sérsic fitting of local bars often includes an additional parameter, $r_{\mathrm{bar}}$, the major-axis radius beyond which the profile is truncated to zero (Kim14). At $z\sim 1.5$, however, we cannot include $r_{\mathrm{bar}}$ as a variable due to the combined effects of low signal-to-noise and flux confusion with other non-axisymmetric disk structures (discussed in BK25). For the same reason, we do not adopt the modified Ferrers profile (laurikainen05), which has a variable $\alpha$ describing the sharpness of the truncation. We plan to revisit this fitting procedure in future work with improved treatment of disk substructures.
Instead, we use bar lengths ($R_{\mathrm{bar}}$) from the ellipse-fit method: the SMA corresponding to the peak ellipticity. This follows the approach of kim21 for galaxies at $z\lesssim 0.84$. liang24 showed that this measure can overestimate bar length when the intrinsic length is << PSF FWHM. However, detectability also drops sharply below this threshold. Since all bars in our sample have SMA values $\gtrsim$ PSF FWHM, we assume they are only detected because their intrinsic lengths are also $\gtrsim$ PSF FWHM. We therefore adopt an uncertainty of $10\%$ (as measured by liang24, for bars $\gtrsim$ PSF FWHM). We note, however, that the uncertainties in liang24 were derived from local galaxies after applying redshift-dependent systematic effects and disk scaling relations; intrinsic (sub-)structural properties at high redshift (guo12; guo15; sattari23; kalita24; kalita25b; mercier25) were not considered. Investigating this effect is beyond the scope of this work. Finally, we correct for the projection effects, by using the axis ratio of the disk ($q$) and the position angle difference between the major axes of the disk and the bar ($\theta$):
@@ 3.5 The bar length uncertainty and its implication
The lack of truncation in our model may still bias some measurements, especially the bar Sérsic index. If truncation is present in the data but not modeled, the Sérsic index may be artificially lowered to account for the flux drop. To test this, we performed a second round of 2D fitting with a fixed-radius truncation, using the ellipse-derived bar length. The ellipse corresponding to the bar (Sec.3.1) was used as the bar edge, beyond which the bar profile was masked. We then re-ran the MCMC sampling, starting from the previous best-fit values. All bar profile measurements, particularly the Sérsic index, remained consistent with the original values within their uncertainties. We therefore conclude that the flatness of the bar profiles indicated by the Sérsic index is robust, and consistent with the flat profiles observed both in the visual 1D major-axis profiles and the 1D fits (Sec.3.3).
4 Results
@@
4.1 Bar profile
The first key property we study is the bar major-axis profile, which spans a range from flat to exponential. Using the 1D flux distribution along the bar ellipse major axis (Sec. 3.1), we identify clear shoulders in 5/9 galaxies (Figs. 5, 3), suggestive of already-present flat bars. This qualitative assessment is further supported by the 1D and 2D Sérsic fits. For comparison with the literature, we adopt the classification scheme of kim15, which divides bars into flat ($n<0.4$), intermediate ($0.4<n<0.8$), and exponential ($n>0.8$). This highlights that bar surface brightness profiles are not strictly flat or exponential, but rather span a continuum.
Figure 6: Bar profiles in 1D and 2D: Sérsic index measurements for each galaxy in our sample. The x-axis shows the 1D fitting results (Sec.3.3), while the y-axis shows the 2D fitting results (Sec.3.4). The shaded regions indicate the classification scheme: flat ($n<0.4$), intermediate ($0.4<n<0.8$), and exponential ($n>0.8$).
The 1D fits indicate that eight bars in our sample fall in the flat-to-intermediate regime ($n<0.8$), while the 1D fitting failed in the final one (ID: 852; albeit it clearly shows shoulders in the major-axis profile). However, the major-axis profile does not capture the full 2D flux distribution, and bar models may still be influenced by residual disk substructures, even after masking the primary spiral arms (Sec. 3.2). This is evident in some fits where the composite model converges to a nonphysical configuration, with the bar fitting large-scale outer-disk fluctuations. To mitigate this, we constrain the effective radius to $\lesssim 1/2\times$ the bar ellipse major-axis length. After applying this constraint, the 1D Sérsic indices remain consistent with the visual major-axis profiles, with bars showing clear shoulders also exhibiting low Sérsic indices.
@@ 4.1 Bar profile
4.2 Bar lengths
Figure 7: $R_{\mathrm{bar}}$ vs. galaxy stellar mass. Also shown are values for late- and early-type barred galaxies at $0.2<z<0.85$ from kim21, along with their median and $1\,\sigma$ scatter.
Figure 8: $R_{\mathrm{bar}}$ normalized by the host disk effective radius vs. galaxy stellar mass for our sample. Gray points show galaxies at $0.5<z<0.85$ from kim21. Their median and $1\,\sigma$ scatter are shown in red.
Figure 9: Bar Sérsic index vs. $R_{\mathrm{bar}}$ for our sample. Values for massive ($M_{\mathrm{*}}>10^{10}\,\rm M_{\mathrm{\odot}}$) local galaxies from kim15 are also shown for comparison.
@@ 4.2 Bar lengths
The bar lengths ($R_{\mathrm{bar}}$) measured from ellipse fitting allow us to compare with bars at $z<1$. As shown in Fig.7, our bars lie well below the median of the distribution observed at $0.2<z<0.85$ by kim21, at $1\,\sigma$ of their scatter. A lower offset is seen in the normalized bar lengths (with the disk $r_{e}$ being the normalizing factor; Fig.8). Half of our values are within $\pm 1\,\sigma$ scatter around the median of their distribution, while the rest are below it. In both cases, we cannot determine any correlation with stellar mass from our data (as expected in the local Universe; erwin18; kim21), due to the small sample size.
In Fig. 9, the two exponential bars are shorter by $\sim 1-2\,\rm kpc$ compared to the mean of the flatter bars in our sample. This trend between bar length and Sérsic index agrees with the distribution of $10^{10}\,M_{\mathrm{*}}/M_{\mathrm{\odot}}$ galaxies at $0.2<z<0.85$ in kim21. This likely suggests that mature bars are longer, since the flattening of bars due to trapping stellar orbits also results in an increase in its length. However, if we plot the ratio of the effective radii ($R_{\mathrm{e}}$) of bar to disk(Fig. 10), we find no clear correlation.
4.3 Bar flux ratio and axis ratio
Theoretical studies conclude that bar growth is driven by the loss of angular momentum, which traps disk particles (stars) in the bar potential (debattista00; athanassoula02; athanassoular03; valenzuela03; collier19a; collier19b; ansar24). In a more detailed sense, as the elongated bar (stellar) structure rotates with a constant angular velocity within a differentially rotating disk, additional stars in the disk feel its gravitational effect. Below the co-rotation radius (the radius at which the linear velocity of the bar and the disk match), the stars are moving faster than the bar. Some of these would lose angular momentum to the bar and align themselves with the elongated bar orbits. This reinforces the bar, thereby strengthening it. It should be noted however, that this process occurs primarily within the co-rotation radius and hence the bar length is lower than it (aguerri98). Gradually the bar slows down due to a net loss of angular momentum, which is expelled to the outer disk and the surrounding halo. As this happens, the co-rotation radius also increases, as the linear velocities of the bar and the disk will now match at a higher radius. Hence, the bar is allowed to get longer, with greater ellipticity (athanassoula02b), as more stars align themselves with the bar.
Figure 11: The distribution of $S_{\mathrm{bar}}$ (defined in Eq.3) and the axis ratio ($\equiv 1-\epsilon$), both measured from the 2D fitting in Sec. 3.4. We find a correlation coefficient $r=-0.42\pm 0.18$.
@@ 4.3 Bar flux ratio and axis ratio
We choose this model-based parametrization rather than the classical $A_{\mathrm{2}}$ (the Fourier amplitude of the $m=2$ component relative to $m=0$; ohta90; laurikainen02; diazgarcia16; lee19) for three reasons. First, we avoid biases from imperfect deprojection: if the disk is not fully face-on, the $m=2$ amplitude is artificially boosted. Second, bulges are not always spherical, so even a correct deprojection can induce additional non-axisymmetries. This problem is amplified because the central regions of these $z\sim 1.5$ galaxies are only marginally resolved. Finally, Cosmic Noon galaxies contain substantial substructures such as clumps and spirals (kalita25a; kalita24c), which also contribute to the $m=2$ signal.
4.4 Quantifying the effect of bars
Measurement of the relative flux contained within the bar (e.g., $S_{\mathrm{bar}}$) provides only a photometric estimate of bar strength, and does not directly quantify the physical effect of the bar on its host galaxy. However, seidel15 found that spectroscopy-based kinematic torques induced by bars are well correlated with the maximum transverse-to-radial force ratio ($Q_{\mathrm{b}}$; combes81; buta01; laurikainen02; buta04; diazgarcia16; lee20), which can be calculated directly from galaxy images. To date, $Q_{\mathrm{b}}$ has only been measured at $z<0.8$ (kim21). At higher redshifts, low signal-to-noise ratios and the increasingly complex morphologies of galaxies, make such measurements challenging.
@@ 4.4 Quantifying the effect of bars
At $z>1$, any $Q_{\mathrm{b}}$ estimate would reflect the net effect of all non-axisymmetric substructures, since disks at these epochs host abundant clumps and spirals (elmegreen05; elmegreen07; forster11; guo18; kalita24; kalita25b; kalita25a; kalita25c; mercier25; salcedo25). Isolating the bar’s contribution is therefore challenging. To address this, we adopt a modified $Q_{\mathrm{b}}$ measurement that uses our 2D bulge–disk–bar models (Sec. 3.4) instead of the direct galaxy images. This removes the contribution of other non-axisymmetric features. For the original procedure, we refer the readers to lee20.
We first deproject the model image using the disk $e_{\mathrm{1}}$ and $e_{\mathrm{2}}$ values (which jointly give the axis ratio and the position angle), yielding a face-on surface brightness profile. This is converted to a mass surface density, $\Sigma(x,y)$, assuming a constant mass-to-light ratio derived from the total F [^4] W flux and stellar mass (kalita25a). This approximation may still result in radial variations since the central regions maybe highly obscured. However, our choice of the longest band limits this effect. Moreover, we opt to not use a stellar mass map derived from a pixel-by-pixel SED fitting, since the JWST data only has 4 wavelength bands. We therefore find artificial fluctuations of the order of $\sim 10\%$ of the average value in the mass map, introduced due to fitting uncertainties. Given that our analysis depends on the local gradients, it is highly susceptible to such fluctuations.
@@
where $h_{z}$ is the scale height. Since our galaxies are nearly face-on (axis ratio >0.6>0.6; BK25), we adopt $h_{z}=0.8\pm 0.3\,\rm kpc$, which is twice the scale height determined by lian24 using a JWST-based statistical study of edge-on galaxies. The factor of two is added since the $sech^{2}$ function used in their study drops twice as fast as the exponential function used in this work. The thickness factor, when integrated over the full vertical extent, reduces to unity. This is essential as we are using the integrated $\Sigma(x,y)$ for the whole disk, rather than a surface density for each thin disk component we assume the disk is made up of. Therefore the final density of the disk is given as:
@@
5 Discussion
5.1 The presence of “mature” bars
Our work has leveraged the high-resolution imaging of JWST/NIRCam to determine the major-axis profiles of bars. Using both the 1D (Sec.3.3) and 2D (Sec.3.4) analyses, we find that seven of the nine barred galaxies have flat-to-intermediate ($n<0.8$) profiles, while the other two lie in the exponential regime ($n>0.8$). The implications of non-exponential profiles has been debated since the pioneering study of elmegreen85. These profiles, generally referred to as flat bars (and more recently subdivided into flat and intermediate by kim15), are mainly found in early-type systems in the local Universe (elmegreen85; buamgart86; elmegreen96c; regan97). This bias has been interpreted as evidence that flat bars evolve from exponential bars, growing longer and stronger in the process (elmegreen89; kim15; lee19). More recently, erwin23 showed that bar profile correlates more fundamentally with galaxy mass rather than Hubble type, while kruk18 suggested an evolutionary trend based on barred galaxy disks being redder, and hence lower in star-formation, than disks in unbarred galaxies. Direct evidence for profile evolution has also been found in N-body simulations: anderson22 showed that bars begin with exponential profiles and evolve into flat profiles in $\sim 1-2\,$Gyr.
Figure 13: Distribution of $Q_{\mathrm{b}}$ as a function of bar axis ratio, divided into three profile categories. Also shown is the lee20 distribution of local barred galaxies. The grey region denotes $Q_{\mathrm{b}}$ values $\leq 0.1$, the uncertainty threshold signifying the maximum systematic fluctuations of the measurement procedure.

### v2

@@ III.2 Masking the spiral arms
III.5 The bar length uncertainty and its implication
We fit the bars in our sample galaxies with a Sérsic profile. However, this does not allow us to measure bar length. Physically, the bar length is the maximum extent of the $x_{\mathrm{1}}$ orbits of the trapped stars making up the bar (J. Binney & S. Tremaine, 2008). In the local Universe, this typically corresponds to the major axis derived from ellipse fits to surface brightness isophotes (Sec. III.1), beyond which the bar surface brightness drops rapidly. For this reason, 2D Sérsic fitting of local bars often includes an additional parameter, $r_{\mathrm{bar}}$, the major-axis radius beyond which the profile is truncated to zero (T. Kim et al., 2014). At $z\sim 1.5$, however, we cannot include $r_{\mathrm{bar}}$ as a variable due to the combined effects of low signal-to-noise and flux confusion with other non-axisymmetric disk structures (discussed in BK25). For the same reason, we do not adopt the modified Ferrers profile (E. Laurikainen et al., 2005), which has a variable $\alpha$ describing the sharpness of the truncation. We plan to revisit this fitting procedure in future work with improved treatment of disk substructures.
Instead, we use bar lengths ($R_{\mathrm{bar}}$) from the ellipse-fit method: the SMA corresponding to the peak ellipticity. This follows the approach of T. Kim et al. (2021) for galaxies at $z\lesssim 0.84$. X. Liang et al. (2024) showed that this measure can overestimate bar length when the intrinsic length is << PSF FWHM. However, detectability also drops sharply below this threshold. Since all bars in our sample have SMA values $\gtrsim$ PSF FWHM, we assume they are only detected because their intrinsic lengths are also $\gtrsim$ PSF FWHM. We therefore adopt an uncertainty of $10\%$ (as measured by X. Liang et al., 2024, for bars $\gtrsim$ PSF FWHM). We note, however, that the uncertainties in X. Liang et al. (2024) were derived from local galaxies after applying redshift-dependent systematic effects and disk scaling relations; intrinsic (sub-)structural properties at high redshift (Y. Guo et al., 2012, 2015; Z. Sattari et al., 2023; B. S. Kalita et al., 2024a, 2025a; W. Mercier et al., 2025) were not considered. Investigating this effect is beyond the scope of this work. Finally, we correct for the projection effects, by using the axis ratio of the disk ($q$) and the position angle difference between the major axes of the disk and the bar ($\theta$):
@@ III.5 The bar length uncertainty and its implication
The lack of truncation in our model may still bias some measurements, especially the bar Sérsic index. If truncation is present in the data but not modeled, the Sérsic index may be artificially lowered to account for the flux drop. To test this, we performed a second round of 2D fitting with a fixed-radius truncation, using the ellipse-derived bar length. The ellipse corresponding to the bar (Sec.III.1) was used as the bar edge, beyond which the bar profile was masked. We then re-ran the MCMC sampling, starting from the previous best-fit values. All bar profile measurements, particularly the Sérsic index, remained consistent with the original values within their uncertainties. We therefore conclude that the flatness of the bar profiles indicated by the Sérsic index is robust, and consistent with the flat profiles observed both in the visual 1D major-axis profiles and the 1D fits (Sec.III.3).
IV Results
The final compilation of the bar properties of the galaxies in our sample is provided in Table 1. We now discuss these parameters in the following subsections.
@@ IV Results
IV.1 Bar profile
The first key property we study is the bar major-axis profile, which spans a range from flat to exponential. Using the 1D flux distribution along the bar ellipse major axis (Sec. III.1), we identify clear shoulders in 5/9 galaxies (Figs. 5, 3), suggestive of already-present flat bars. This qualitative assessment is further supported by the 1D and 2D Sérsic fits. For comparison with the literature, we adopt the classification scheme of T. Kim et al. (2015), which divides bars into flat ($n<0.4$), intermediate ($0.4<n<0.8$), and exponential ($n>0.8$). This highlights that bar surface brightness profiles are not strictly flat or exponential, but rather span a continuum.
Figure 6: Bar profiles in 1D and 2D: Sérsic index measurements for each galaxy in our sample. The x-axis shows the 1D fitting results (Sec.III.3), while the y-axis shows the 2D fitting results (Sec.III.4). The shaded regions indicate the classification scheme: flat ($n<0.4$), intermediate ($0.4<n<0.8$), and exponential ($n>0.8$).
The 1D fits indicate that eight bars in our sample fall in the flat-to-intermediate regime ($n<0.8$), while the 1D fitting failed in the final one (ID: 852; albeit it clearly shows shoulders in the major-axis profile). However, the major-axis profile does not capture the full 2D flux distribution, and bar models may still be influenced by residual disk substructures, even after masking the primary spiral arms (Sec. III.2). This is evident in some fits where the composite model converges to a nonphysical configuration, with the bar fitting large-scale outer-disk fluctuations. To mitigate this, we constrain the effective radius to $\lesssim 1/2\times$ the bar ellipse major-axis length. After applying this constraint, the 1D Sérsic indices remain consistent with the visual major-axis profiles, with bars showing clear shoulders also exhibiting low Sérsic indices.
@@ IV.1 Bar profile
IV.2 Bar lengths
Figure 7: $R_{\mathrm{bar}}$ vs. galaxy stellar mass. Also shown are values for late- and early-type barred galaxies at $0.2<z<0.85$ from T. Kim et al. (2021), along with their median and $1\,\sigma$ scatter.
Figure 8: $R_{\mathrm{bar}}$ normalized by the host disk effective radius vs. galaxy stellar mass for our sample. Gray points show galaxies at $0.5<z<0.85$ from T. Kim et al. (2021). Their median and $1\,\sigma$ scatter are shown in red.
Figure 9: Bar Sérsic index vs. $R_{\mathrm{bar}}$ for our sample. Values for massive ($M_{\mathrm{*}}>10^{10}\,\rm M_{\mathrm{\odot}}$) local galaxies from T. Kim et al. (2015) are also shown for comparison.
@@ IV.2 Bar lengths
The bar lengths ($R_{\mathrm{bar}}$) measured from ellipse fitting allow us to compare with bars at $z<1$. As shown in Fig.7, our bars lie well below the median of the distribution observed at $0.2<z<0.85$ by T. Kim et al. (2021), at $1\,\sigma$ of their scatter. A lower offset is seen in the normalized bar lengths (with the disk $R_{e}$ being the normalizing factor; Fig.8). Half of our values are within $\pm 1\,\sigma$ scatter around the median of their distribution, while the rest are below it. In both cases, we cannot determine any correlation with stellar mass from our data (as expected in the local Universe; P. Erwin, 2018; T. Kim et al., 2021), due to the small sample size.
In Fig. 9, the two exponential bars are shorter by $\sim 1-2\,\rm kpc$ compared to the mean of the flatter bars in our sample. This trend between bar length and Sérsic index agrees with the distribution of $10^{10}\,M_{\mathrm{*}}/M_{\mathrm{\odot}}$ galaxies at $0.2<z<0.85$ in T. Kim et al. (2021). This likely suggests that mature bars are longer, since the flattening of bars due to trapping stellar orbits also results in an increase in its length. However, if we plot the ratio of the effective radii ($R_{\mathrm{e}}$) of bar to disk(Fig. 10), we find no clear correlation.
IV.3 Bar flux ratio and axis ratio
Theoretical studies conclude that bar growth is driven by the loss of angular momentum, which traps disk particles (stars) in the bar potential (V. P. Debattista & J. A. Sellwood, 2000; E. Athanassoula, 2002, 2003; O. Valenzuela & A. Klypin, 2003; A. Collier et al., 2019a, b; S. Ansar & M. Das, 2024). In a more detailed sense, as the elongated bar (stellar) structure rotates with a constant angular velocity within a differentially rotating disk, additional stars in the disk feel its gravitational effect. Below the co-rotation radius (the radius at which the linear velocity of the bar and the disk match), the stars are moving faster than the bar. Some of these would lose angular momentum to the bar and align themselves with the elongated bar orbits. This reinforces the bar, thereby strengthening it. It should be noted however, that this process occurs primarily within the co-rotation radius and hence the bar length is lower than it (J. A. L. Aguerri et al., 1998). Gradually the bar slows down due to a net loss of angular momentum, which is expelled to the outer disk and the surrounding halo. As this happens, the co-rotation radius also increases, as the linear velocities of the bar and the disk will now match at a higher radius. Hence, the bar is allowed to get longer, with greater ellipticity (E. Athanassoula & A. Misiriotis, 2002), as more stars align themselves with the bar.
Figure 11: The distribution of $S_{\mathrm{bar}}$ (defined in Eq.3) and the axis ratio ($\equiv 1-\epsilon$), both measured from the 2D fitting in Sec. III.4. We find a correlation coefficient $r=-0.42\pm 0.18$.
@@ IV.3 Bar flux ratio and axis ratio
We choose this model-based parametrization rather than the classical $A_{\mathrm{2}}$ (the Fourier amplitude of the $m=2$ component relative to $m=0$; K. Ohta et al., 1990; E. Laurikainen & H. Salo, 2002; S. Díaz-García et al., 2016; Y. H. Lee et al., 2019) for three reasons. First, we avoid biases from imperfect deprojection: if the disk is not fully face-on, the $m=2$ amplitude is artificially boosted. Second, bulges are not always spherical, so even a correct deprojection can induce additional non-axisymmetries. This problem is amplified because the central regions of these $z\sim 1.5$ galaxies are only marginally resolved. Finally, Cosmic Noon galaxies contain substantial substructures such as clumps and spirals (B. S. Kalita et al., 2025b, 2024b), which also contribute to the $m=2$ signal.
IV.4 Quantifying the effect of bars
Measurement of the relative flux contained within the bar (e.g., $S_{\mathrm{bar}}$) provides only a photometric estimate of bar strength, and does not directly quantify the physical effect of the bar on its host galaxy. However, M. K. Seidel et al. (2015) found that spectroscopy-based kinematic torques induced by bars are well correlated with the maximum transverse-to-radial force ratio ($Q_{\mathrm{b}}$; F. Combes & R. H. Sanders, 1981; R. Buta & D. L. Block, 2001; E. Laurikainen & H. Salo, 2002; R. Buta et al., 2004; S. Díaz-García et al., 2016; Y. H. Lee et al., 2020), which can be calculated directly from galaxy images. To date, $Q_{\mathrm{b}}$ has only been measured at $z<0.8$ (T. Kim et al., 2021). At higher redshifts, low signal-to-noise ratios and the increasingly complex morphologies of galaxies, make such measurements challenging.
@@ IV.4 Quantifying the effect of bars
At $z>1$, any $Q_{\mathrm{b}}$ estimate would reflect the net effect of all non-axisymmetric substructures, since disks at these epochs host abundant clumps and spirals (B. G. Elmegreen & D. M. Elmegreen, 2005; D. M. Elmegreen et al., 2007; N. M. Förster Schreiber et al., 2011; Y. Guo et al., 2018; B. S. Kalita et al., 2024a, 2025a, 2025b, 2025c; W. Mercier et al., 2025; J. M. Espejo Salcedo et al., 2025). Isolating the bar’s contribution is therefore challenging. To address this, we adopt a modified $Q_{\mathrm{b}}$ measurement that uses our 2D bulge–disk–bar models (Sec. III.4) instead of the direct galaxy images. This removes the contribution of other non-axisymmetric features. For the original procedure, we refer the readers to Y. H. Lee et al. (2020).
We first deproject the model image using the disk $e_{\mathrm{1}}$ and $e_{\mathrm{2}}$ values (which jointly give the axis ratio and the position angle), yielding a face-on surface brightness profile. This is converted to a mass surface density, $\Sigma(x,y)$, assuming a constant mass-to-light ratio derived from the total F [^4] W flux and stellar mass (B. S. Kalita et al., 2025b). This approximation may still result in radial variations since the central regions maybe highly obscured. However, our choice of the longest band limits this effect. Moreover, we opt to not use a stellar mass map derived from a pixel-by-pixel SED fitting, since the JWST data only has 4 wavelength bands. We therefore find artificial fluctuations of the order of $\sim 10\%$ of the average value in the mass map, introduced due to fitting uncertainties. Given that our analysis depends on the local gradients, it is highly susceptible to such fluctuations.
@@
where $h_{z}$ is the scale height. Since our galaxies are nearly face-on (axis ratio >0.6>0.6; BK25), we adopt $h_{z}=0.8\pm 0.3\,\rm kpc$, which is twice the scale height determined by J. Lian & L. Luo (2024) using a JWST-based statistical study of edge-on galaxies. The factor of two is added since the $sech^{2}$ function used in their study drops twice as fast as the exponential function used in this work. The thickness factor, when integrated over the full vertical extent, reduces to unity. This is essential as we are using the integrated $\Sigma(x,y)$ for the whole disk, rather than a surface density for each thin disk component we assume the disk is made up of. Therefore the final density of the disk is given as:
@@
V Discussion
As we have been able to characterize the morphological structure of bars in our sample at $z\sim 1.5$, we will now discuss the implications.
V.1 The presence of “mature” bars
Our work has leveraged the high-resolution imaging of JWST/NIRCam to determine the major-axis profiles of bars. Using both the 1D (Sec.III.3) and 2D (Sec.III.4) analyses, we find that seven of the nine barred galaxies have flat-to-intermediate ($n<0.8$) profiles, while the other two lie in the exponential regime ($n>0.8$). The implications of non-exponential profiles has been debated since the pioneering study of B. G. Elmegreen & D. M. Elmegreen (1985). These profiles, generally referred to as flat bars (and more recently subdivided into flat and intermediate by T. Kim et al., 2015), are mainly found in early-type systems in the local Universe (B. G. Elmegreen & D. M. Elmegreen, 1985; C. W. Baumgart & C. J. Peterson, 1986; B. G. Elmegreen et al., 1996; M. W. Regan & D. M. Elmegreen, 1997). This bias has been interpreted as evidence that flat bars evolve from exponential bars, growing longer and stronger in the process (B. G. Elmegreen & D. M. Elmegreen, 1989; T. Kim et al., 2015; Y. H. Lee et al., 2019). More recently, P. Erwin et al. (2023) showed that bar profile correlates more fundamentally with galaxy mass rather than Hubble type, while S. J. Kruk et al. (2018) suggested an evolutionary trend based on barred galaxy disks being redder, and hence lower in star-formation, than disks in unbarred galaxies. Direct evidence for profile evolution has also been found in N-body simulations: S. R. Anderson et al. (2022) showed that bars begin with exponential profiles and evolve into flat profiles in $\sim 1-2\,$Gyr.
Figure 13: Distribution of $Q_{\mathrm{b}}$ as a function of bar axis ratio, divided into three profile categories. Also shown is the Y. H. Lee et al. (2020) distribution of local barred galaxies. The grey region denotes $Q_{\mathrm{b}}$ values $\leq 0.1$, the uncertainty threshold signifying the maximum systematic fluctuations of the measurement procedure.